OpenStudy (jake4272):

Write the equation for 5x + 2y = 3 in slope-intercept form. y =5/2 x - 3/5 y = -5/2x + 3/2 y = -2/5x + 3/5

sam (.sam.):

What's your progress? Any ideas?

OpenStudy (jake4272):

besides what my teachers have always told me if you dont know it guess b

OpenStudy (prathamesh_m):

You basically need to get y on one side. So simply imagine it to be a linear equation in y and solve for y.

OpenStudy (jake4272):

im slow i really dont know these kind of things

OpenStudy (prathamesh_m):

Ok do you know how to isolate y in that equation?

OpenStudy (prathamesh_m):

Start by subtracting 5x on both sides.

OpenStudy (jake4272):

kinda sorta

OpenStudy (jake4272):

ok

OpenStudy (shamim):

Required equation is like y=mx+c

OpenStudy (shamim):

m=slope of line

OpenStudy (shamim):

c=y intercept

OpenStudy (jake4272):

ok i see where this going continue

OpenStudy (shamim):

U hv 5x+2y=3 2y=3-5x y=?

OpenStudy (shamim):

Just divide both sides by 2

OpenStudy (jake4272):

ok

OpenStudy (shamim):

Wt did u get